OpenStudy (anonymous):

An object in motion has Kinetic energy, with 2000N of force over 1000M, Total energy = 2,000,000J If that object is moving with that energy, suddenly increased the force to x10 times of its original =20,000N The total energy of the object now is: 20,000,000J

OpenStudy (anonymous):

Generally if force is increased in the system or the distance that the force is being applied. That's change in the kinetic energy right? W=Fd Initial W= 2,000,000j After increase of F, W= 20,000,000j I am I doing it right here? @Jemurray3 @Vincent-Lyon.Fr @experimentX

OpenStudy (anonymous):

Lets say we had to systems. System 1- D = 10M, F= 1000N System 2- Same D= 10M F=100N Both system are similar in every way. But only the force is different. That would mean the energy of system 2 is less right?

OpenStudy (anonymous):

@Jemurray3 Based on what we've talked about before I "THINK" im still on the right track here.

OpenStudy (anonymous):

@Jemurray3 @Vincent-Lyon.Fr @experimentX I forgot a lot of material. The answer to my question is yes. By increasing the applied force in the system you will increase the total energy in the system. In whatever form it is, by increasing on vector that is important to give you the total energy it would eventually increase the total energy of the system. I answered myself I guess. :)

OpenStudy (anonymous):

@Visionary01 By looking at W= Fd you'd figure the answer to your question :P In any system where work is being done... By increase d or F you'd eventually increase the system energy so you're right so far :) I don't find anything wrong here.

OpenStudy (anonymous):

Thanks guys for the medal's :)

OpenStudy (anonymous):

If you have applied 2000N over 1000m and now you change the force to 20000N your energy remains 2,000,000 J until you move the object with the new force whereupon the energy increases from 2,000,000 J to 2,000,000J + 20000N*the number of meters you move with the new force. so after 1m you now have 2,020,000 J after 2 m 2,040,000 J and so on. KE= F1d1+F2d2+F3d3....

OpenStudy (anonymous):

HUGE confusion here! right! If the force of the first system increased over the same distance which is 1000M Yea, you'd get a higher increase. @gleem you're adding up the force over 1m that wrong. Its the same distance being applied. Generally, when a system increases its input force and has the SAME distance the total energy will be higher than before.

OpenStudy (anonymous):

The drag coefficient is a number , a proportionality "constant" , that relate the drag ie, resistance force of a fluid flowing around a moving object to its speed, the density of the fluid and an effective area related to the projected area of the object facing the flow and its shape. .\[ F _{drag}=c _{d}*\left( \left( \frac{ 1 }{ 2} \right)\rho v ^{2} A _{eff}\right)\] The drag coefficient is not really a constant, but varies in a complex way as a function of speed, flow direction, object position, object size, fluid density and fluid viscosity.